OpenStudy (anonymous):

an architect designs a rectangle flower garden such that the width is exactly 2/3 of the length if 280 feet of antique picket fencing are to be used to enclosure the garden find the dimension of the garden.

OpenStudy (anonymous):

84*56 are the dimensions, with 84 being length and width equaling 56.

OpenStudy (anonymous):

thanks !!! what about V=BQS for B

OpenStudy (anonymous):

What are you talking about there? Solve for B? Then B = V / QS when Q*S does not equal 0.

OpenStudy (anonymous):

thats what the problem say

OpenStudy (anonymous):

Then what I typed B = V / QS when Q*S does not equal 0.

OpenStudy (anonymous):

ok what bout V=ARS for A

OpenStudy (anonymous):

nvm I got that what is 3x-5y=19 for y

OpenStudy (anonymous):

\[y=\frac{ 3x }{ 5 }*\frac{ 19 }{5}\]

OpenStudy (anonymous):

Btw, mark best response or something...

OpenStudy (anonymous):

the perimeter of a basketball court is 114 meters and the length is 6 meters long then twice the length what are the length and width

OpenStudy (anonymous):

40m L, 17m W.

OpenStudy (anonymous):

thanks so much im bad at math

OpenStudy (anonymous):

Find the area and perimeter of the room (neglecting doors)identify whether baseboard has to do with area or perimeter and the same with carpet 9ft and 7 ft